# Towards Extreme Image Compression with Latent Feature Guidance and Diffusion Prior

### v3

@@ Abstract
Image compression at extremely low bitrates (below 0.1 bits per pixel (bpp)) is a significant challenge due to substantial information loss. In this work, we propose a novel two-stage extreme image compression framework that exploits the powerful generative capability of pre-trained diffusion models to achieve realistic image reconstruction at extremely low bitrates. In the first stage, we treat the latent representation of images in the diffusion space as guidance, employing a VAE-based compression approach to compress images and initially decode the compressed information into content variables. The second stage leverages pre-trained stable diffusion to reconstruct images under the guidance of content variables. Specifically, we introduce a small control module to inject content information while keeping the stable diffusion model fixed to maintain its generative capability. Furthermore, we design a space alignment loss to force the content variables to align with the diffusion space and provide the necessary constraints for optimization. Extensive experiments demonstrate that our method significantly outperforms state-of-the-art approaches in terms of visual performance at extremely low bitrates.
@@ I Introduction
Recently, diffusion models have exhibited impressive generation ability in image and video generation [13, 14, 15], encouraging researchers to develop various diffusion-based perception-driven compression methods [16, 17, 18, 19]. For extreme image compression, some works leverage pre-trained text-to-image diffusion models as prior knowledge to achieve realistic reconstructions at extremely low bitrates. For instance, Pan et al. [20] encode images as textual embeddings with extremely low bitrates, using pre-trained text-to-image diffusion models for realistic reconstruction. Lei et al. [21] directly transmit short text prompts and compressed image sketches, employing the pre-trained ControlNet [14] to produce reconstructions with high perceptual quality and semantic fidelity. However, these methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capabilities of pre-trained diffusion models, resulting in reconstruction results that are inconsistent with the original image (see Fig. 1(d)). Therefore, how to develop an effective diffusion-based extreme generative compression method is worth further exploration.
@@ II-B Extreme Image Compression
In addition, some approaches use prior knowledge to achieve extreme image compression. Yue et al. [36] describe input images based on the down-sampled version and handcrafted features, and use these descriptions to reconstuct the images from a large-scale image database. Their method can achieve impressive compression performance when the large-scale image database contains images that are highly correlated with the input images. Benefiting from the bijective and information-lossless property of invertible neural networks (INNs), Gao et al. leverage INNs to mitigate the significant information loss in extreme image compression [37]. Wei et al. employ invertible and generative priors to achieve extreme compression by rescaling images with extreme scaling factors (i.e., 16$\times$ and 32$\times$) [38]. In [39], Li et al. employ physical priors (i.e., attenuation coefficient and ambient light) and the semantic prior for extreme underwater image compression, which may not generalize well to natural images with different scenarios. Jiang et al. utilize text descriptions as prior to guide image compression for better compression performance [40].
Inspired by the tremendous success of diffusion models in image generation, some methods [20, 21] use more powerful pre-trained text-to-image diffusion models as prior knowledge. In [20], Pan et al. encode images into short text embeddings and then generate high-quality images with pre-trained text-to-image diffusion models by feeding the text embeddings. Lei et al. [21] directly compress the short text prompts and binary contour sketches on the encoded side, and then use them as input to the pre-trained text-to-image diffusion model for reconstruction on the decoded side. However, these diffusion-based methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capabilities of pre-trained diffusion models.
@@ III-A2 Decoding process
Given the quantized $\hat{y}$ and $\hat{z}$, we first use the information extraction network $f_{c}$ to extract a representation $w$ from $\hat{z}$, which can be expressed as:
@@ III-A3 Network Details
Fig. 3 illustrates the network architecture of LFGCM. The information extraction network $f_{c}$ has the same network structure as the hyper-decoder $\mathcal{N}_{hd}$, and we adopt context model $\mathcal{C}_{m}$ proposed by He et al. [31]. The guidance components (the elements denoted by red arrows, $SFT$ and $SFT\ Resblk$) first use a series of convolutions to resize the external feature $G$ to the appropriate dimensions. Then the SFT layers [45] are employed to inject the network with external guidance information. Specifically, given an external feature $G$ and an intermediate feature map $F$, a pair of affine transformation parameters (i.e., $\alpha$ for scaling and $\beta$ for shifting) is generated as follows:
@@ III-C3 Space Alignment Loss
where $\lambda_{sa}$ and $\lambda_{ne}$ denote the weights for space alignment loss and noise estimation loss, respectively. $\lambda$ is used to achieve a trade-off between rate and distortion.
@@ IV-A1 Implementation
We train DiffEIC on the LSDIR [50] dataset, which contains 84,991 high-quality training images. The images are randomly cropped to 512$\times$512 resolution. In our experiments, we use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the diffusion prior. We train our model in an end-to-end manner using Eq. (13), where $\lambda_{sa}$ and $\lambda_{ne}$ are set as 2 and 1, respectively. To achieve different coding bitrates, we choose $\lambda$ from $\{1,2,4,8,16\}$. For optimization, we utilize Adam [51] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and set the learning rate to 1×$\times 10^{-4}$. The training batch size is set to 4. Inspired by previous work [6], we first train the proposed DiffEIC with $\lambda=1$ for 300K iterations, and then adapt them using target $\lambda$ for another 200K iterations. We set the learning rate to 2×$\times 10^{-5}$ during the fine-tuning process. For inference, we adopt spaced DDPM sampling [52] with 50 steps to reconstruct the images. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU. The source code and trained models will be available on the authors’ homepage after acceptance of the manuscript.
@@ IV-B1 Quantitative Comparisons
Fig. 4 shows the rate-perception curves at low bitrates for different methods over the three datasets. It can be observed that the proposed DiffEIC performs much better than BPG, VVC, and ELIC [31] for all perceptual metrics. Although the Text+Sketch [21] achieves the best NIQE value of all the methods, it fails to ensure the pixel fidelity, where the LPIPS value is the highest. For other generative image compression methods, the proposed DiffEIC yields lower DISTS, FID, and KID values, indicating that DiffEIC excels in preserving the perceptual integrity of the images and producing reconstructions with minimal perceptual differences from the originals.
The rate-distortion performance comparison is shown in Fig. 5. Since Text+Sketch [21] has ignored the pixel-level fidelity of the reconstruction results, we do not report its rate-distortion performance. Compared to Wei et al. [38] and PerCo [19], the proposed DiffEIC achieves better PSNR and MS-SSIM values. However, we find that DiffEIC is worse than other comparative methods. The reason behind this is that the proposed DiffEIC uses stable diffusion prior for realistic detail reconstruction at extremely low bitrates, which does not ensure pixel-level accuracy. To further demonstrate this, we report the PSNR and MS-SSIM values of the stable diffusion autoencoder (see the black horizontal line in Fig. 5), which can be treated as the upper bound of the performance of DiffEIC. Although it sacrifices some fidelity, the proposed DiffEIC is able to capture realism at extremely low bitrates.
@@ VI Conclusion
In this paper, we propose a novel extreme image compression framework, named DiffEIC, which combines compressive VAEs with pre-trained text-to-image diffusion models to achieve realistic and high-fidelity reconstructions at extremely low bitrates (below 0.1 bpp). First, we introduce a VAE-based latent feature-guided compression module to adaptively select information essential for reconstruction. This module compresses images and initially decodes them into content variables. The latent feature guidance strategy effectively improves reconstruction fidelity. Second, we propose a conditional diffusion decoding module that leverages the powerful generative capability of pretrained stable diffusion to reconstruct images with realistic details. Finally, we design a simple yet effective space alignment loss to optimize DiffEIC within a unified framework. Extensive experiments demonstrate the superiority of DiffEIC and the effectiveness of the proposed modules.

### v4

@@ Abstract
Image compression at extremely low bitrates (below 0.1 bits per pixel (bpp)) is a significant challenge due to substantial information loss. In this work, we propose a novel two-stage extreme image compression framework that exploits the powerful generative capability of pre-trained diffusion models to achieve realistic image reconstruction at extremely low bitrates. In the first stage, we treat the latent representation of images in the diffusion space as guidance, employing a VAE-based compression approach to compress images and initially decode the compressed information into content variables. The second stage leverages pre-trained stable diffusion to reconstruct images under the guidance of content variables. Specifically, we introduce a small control module to inject content information while keeping the stable diffusion model fixed to maintain its generative capability. Furthermore, we design a space alignment loss to force the content variables to align with the diffusion space and provide the necessary constraints for optimization. Extensive experiments demonstrate that our method significantly outperforms state-of-the-art approaches in terms of visual performance at extremely low bitrates. The source code and trained models are available at https://github.com/huai-chang/DiffEIC.
@@ I Introduction
Recently, diffusion models have exhibited impressive generation ability in image and video generation [13, 14, 15], encouraging researchers to develop various diffusion-based perception-driven compression methods [16, 17, 18, 19]. For extreme image compression, some works leverage pre-trained text-to-image diffusion models as prior knowledge to achieve realistic reconstructions at extremely low bitrates. For instance, Pan et al. [20] encode images as textual embeddings with extremely low bitrates, using pre-trained text-to-image diffusion models for realistic reconstruction. Lei et al. [21] directly transmit short text prompts and compressed image sketches, employing the pre-trained ControlNet [14] to produce reconstructions with high perceptual quality and semantic fidelity. However, these methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capability of pre-trained diffusion models, resulting in reconstruction results that are inconsistent with the original image (see Fig. 1(d)). Therefore, how to develop an effective diffusion-based extreme generative compression method is worth further exploration.
@@ II-B Extreme Image Compression
Some approaches use prior knowledge to achieve extreme image compression. Yue et al. [36] describe input images based on the down-sampled version and handcrafted features, and use these descriptions to reconstuct the images from a large-scale image database. Their method can achieve impressive compression performance when the large-scale image database contains images that are highly correlated with the input images. Benefiting from the bijective and information-lossless property of invertible neural networks (INNs), Gao et al. leverage INNs to mitigate the significant information loss in extreme image compression [37]. Wei et al. employ invertible and generative priors to achieve extreme compression by rescaling images with extreme scaling factors (i.e., 16$\times$ and 32$\times$) [38]. In [39], Li et al. employ physical priors (i.e., attenuation coefficient and ambient light) and the semantic prior for extreme underwater image compression, which may not generalize well to natural images with different scenarios. Jiang et al. utilize text descriptions as prior to guide image compression for better compression performance [40].
Inspired by the tremendous success of diffusion models in image generation, some methods [20, 21] use more powerful pre-trained text-to-image diffusion models as prior knowledge. In [20], Pan et al. encode images into short text embeddings and then generate high-quality images with pre-trained text-to-image diffusion models by feeding the text embeddings. Lei et al. [21] directly compress the short text prompts and binary contour sketches on the encoded side, and then use them as input to the pre-trained text-to-image diffusion model for reconstruction on the decoded side. However, these diffusion-based methods treat pre-trained text-to-image diffusion models as independent components, which limits their ability to fully exploit the generative capability of pre-trained diffusion models.
@@ III-A2 Decoding process
Given the quantized $\hat{y}$ and $\hat{z}$, we first use a information extraction network $f_{c}$ to extract a representation $w$ from $\hat{z}$, which can be expressed as:
@@ III-A3 Network Details
Fig. 3 illustrates the network architecture of LFGCM. The information extraction network $f_{c}$ has the same structure as the hyper-decoder $\mathcal{N}_{hd}$, and we adopt context model $\mathcal{C}_{m}$ proposed by He et al. [31]. The guidance components (the elements denoted by red arrows, $SFT$ and $SFT\ Resblk$) first use a series of convolutions to resize the external feature $G$ to the appropriate dimensions. Then the SFT layers [45] are employed to inject the network with external guidance information. Specifically, given an external feature $G$ and an intermediate feature map $F$, a pair of affine transformation parameters (i.e., $\alpha$ for scaling and $\beta$ for shifting) is generated as follows:
@@ III-C3 Space Alignment Loss
where $\lambda_{sa}$ and $\lambda_{ne}$ denote the weights for space alignment loss and noise estimation loss, respectively. $\lambda$ is used to achieve a trade-off between rate and reconstruction quality.
@@ IV-A1 Implementation
We train DiffEIC on the LSDIR [50] dataset, which contains 84,991 high-quality training images. The images are randomly cropped to 512$\times$512 resolution. In our experiments, we use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the diffusion prior. We train our model in an end-to-end manner using Eq. (13), where $\lambda_{sa}$ and $\lambda_{ne}$ are set to 2 and 1, respectively. To achieve different coding bitrates, we choose $\lambda$ from $\{1,2,4,8,16\}$. For optimization, we utilize Adam [51] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and set the learning rate to 1×$\times 10^{-4}$. The training batch size is set to 4. Inspired by previous work [6], we first train the proposed DiffEIC with $\lambda=1$ for 300K iterations, and then adapt them using target $\lambda$ for another 200K iterations. We set the learning rate to 2×$\times 10^{-5}$ during the fine-tuning process. For inference, we adopt spaced DDPM sampling [52] with 50 steps to reconstruct the images. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
@@ IV-B1 Quantitative Comparisons
Fig. 4 shows the rate-perception curves at low bitrates for different methods over the three datasets. It can be observed that the proposed DiffEIC performs much better than BPG [2], VVC [3], and ELIC [31] for all perceptual metrics. Although the Text+Sketch [21] achieves the best NIQE value of all the methods, it fails to ensure the pixel fidelity, where the LPIPS value is the highest. For other generative image compression methods, the proposed DiffEIC yields lower DISTS, FID, and KID values, indicating that DiffEIC excels in preserving the perceptual integrity of the images and producing reconstructions with minimal perceptual differences from the originals.
The rate-distortion performance comparison is shown in Fig. 5. Since Text+Sketch [21] has ignored the pixel-level fidelity of the reconstruction results, we do not report its rate-distortion performance. Compared to Wei et al. [38] and PerCo [19], the proposed DiffEIC achieves better PSNR and MS-SSIM values. However, we find that DiffEIC is worse than other comparative methods. The reason behind this is that the proposed DiffEIC uses stable diffusion prior for realistic detail reconstruction at extremely low bitrates, which does not ensure pixel-level accuracy. To further demonstrate this, we report the PSNR and MS-SSIM values of the stable diffusion autoencoder (see the black horizontal line in Fig. 5), which can be treated as the upper bound of the performance of DiffEIC. Although it sacrifices some fidelity, the proposed DiffEIC effectively capture realism at extremely low bitrates.
@@ VI Conclusion
In this paper, we propose a novel extreme image compression framework, named DiffEIC, which combines compressive VAEs with pre-trained text-to-image diffusion models to achieve realistic and high-fidelity reconstructions at extremely low bitrates (below 0.1 bpp). First, we introduce a VAE-based latent feature-guided compression module to adaptively select information essential for reconstruction. This module compresses images and initially decodes them into content variables. The latent feature guidance strategy effectively improves reconstruction fidelity. Second, we propose a conditional diffusion decoding module that leverages the powerful generative capability of pre-trained stable diffusion to reconstruct images with realistic details. Finally, we design a simple yet effective space alignment loss to optimize DiffEIC within a unified framework. Extensive experiments demonstrate the superiority of DiffEIC and the effectiveness of the proposed modules.